# The Fréchet Distance Unleashed: Approximating a Dog with a Frog

### v1

@@ 4.4.1 Morphing sensitive simplification
For an $\psi>1$, a Fréchet morphing $\mathcalb{m}$ of $\pi$ and $\sigma$ is $\psi$-approximate morphing if $\mathsf{d}_{\mathcal{F}}\left(\pi,\sigma\right)\leq\omega\left(\mathcalb{m}\right)\leq\psi\mathsf{d}_{\mathcal{F}}\left(\pi,\sigma\right).$ Thus $\mathcalb{l}=\omega\left(\mathcalb{m}\right)/\psi$ is a lower bound on the Fréchet distance between $\pi$ and $\sigma$.
@@
For a morphing $\mathcalb{m}\in\mathcal{M}_{\pi,\sigma}$, a lower bound $\mathcalb{l}\leq\mathsf{d}_{\mathcal{F}}\left(\pi,\sigma\right)$, and a point $p=(x,y)\in\mathcalb{m}$, the slack at $p$ is $\Delta\left(p\right)=\max(\mathcalb{l}-\mathcalb{e}\left(p\right),0)$. A point $p\in\mathcalb{m}$ is right if $\Delta\left(p\right)=0$.
Consider an optimal (say retractable) Fréchet morphing $\mathcalb{m}$ between $\pi$ and $\sigma$. We expect only the point in the morphing realizing the bottleneck to be tight. One can map the slack from the morphing back to the original curves. Specifically, for $v\in\pi\cup\sigma$, let $\max{m}(v)$ be the maximum length leash attached to $v$ during the morphing $\mathcalb{m}$ (this can correspond to the maximum elevation along a vertical or horizontal segment in $\mathcalb{m}$).
@@
For a vertex $v\in\pi\cup\sigma$, the slack of $v$ is $\Delta\left(v\right)=\max(\mathcalb{l}-\max\nolimits{m}(v),0)/4.$
@@ 4.4.2 Computing a lower-bound on the Fréchet distance
$\leq\left\|{\overline{\pi}(\mathcalb{m}_{\pi}(x)\bigr{)}-\overline{\sigma}\bigl{(}\mathcalb{m}_{\sigma}(y)\bigr{)}}\right\|-\left\|{\pi(x)\overline{\pi}(\mathcalb{m}_{\pi}(x)\bigr{)}}\right\|-\left\|{\sigma(y)\overline{\sigma}(\mathcalb{m}_{\sigma}(y)\bigr{)}}\right\|\leq\mathcalb{e}\left(x,y\right).$
@@ 4.4.3 Computing the exact Fréchet distance
Let $\pi$ and $\sigma$ be two curves, and let $\mathcalb{m}$ be a $\psi$-approximate morphing between them, for some $\psi>1$. For $\tau\geq 1$ be some constant, and consider $(\mathcalb{m},\tau)$-sensitive simplifications $\overline{\pi},\overline{\sigma}$, of $\pi,\sigma$, respectively. Let $\mathcalb{m}_{2}:\overline{\pi}\rightarrow\overline{\sigma}$ be the optimal Fréchet morphing between $\overline{\pi}$ and $\overline{\sigma}$, and let $\mathcalb{m}_{1}:\pi\rightarrow\overline{\pi}$, and $\mathcalb{m}_{3}:\overline{\sigma}\rightarrow\sigma$, be the natural morphings associated with the simplifications. This gives rise to a natural morphing $\mathcalb{m}^{\prime}=\mathcalb{m}_{3}\circ\mathcalb{m}_{2}\circ\mathcalb{m}_{1}$ from $\pi$ to $\sigma$. If $\omega\left(\mathcalb{m}^{\prime}\right)={\overline{\mathsf{d}_{\mathcal{F}}}\left(\overline{\pi},\overline{\sigma}\right)}$, then $\mathcalb{m}^{\prime}$ realizes the optimal Fréchet distance between $\pi$ and $\sigma$. where ${\overline{\mathsf{d}_{\mathcal{F}}}\left(\overline{\pi},\overline{\sigma}\right)}$ is the Fréchet distance between $\overline{\pi}$ and $\overline{\sigma}$ under the modified elevation function $\overline{\mathcalb{e}}$.
@@ 4.5.2 The sweep distance
$F(x)=\int f(x)\mathop{\mathrm{d}}\!x=\left(\frac{c}{2\sqrt{a}}-\frac{b^{2}}{8a^{3/2}}\right)\mathop{\mathrm{asinh}}\frac{2ax+b}{\sqrt{4ac-b^{2}}}+\left(\frac{x}{2}+\frac{b}{4a}\right)f(x)$
@@
The previous subsection gives us pricing on the edges of the VE graph (see Definition 3.2). We can then use Dijkstra to compute the shortest path in the VE graph. Naturally, the resulting path is not monotone, but monotonicity can be easily achieved by introducing middle points, as described in Remark 4.6. We then recompute the shortest path getting a morphing, and repeat the refinement step if the morphing is not yet monotone. Usually, one round of refinement seems to suffice. We output the computed distance (for the associated monotone morphing). We refer to this quantity, as the sweep distance between the two original curves, denoted by $\mathsf{d}_{SW}\left(\pi,\sigma\right)$.
@@ 4.5.3 Computing a lower bound on the CDTW distance using the sweep distance
Consider computing the sweep distance for this “elevation” function. Since the sweep distance is decomposed along its $x$ and $y$ components, and the function is a constant inside a grid cell, one can safely assume the optimal morphing is axis aligned. The only remaining possibility is that the optimal morphing enters a column on a left edge, climbs in a stairway, and leave through an edge on the right. Let $e_{1},e_{2},\ldots,e_{k}$ be the horizontal edges this path intersects, and observe that one can always modify this path, to move vertically to the cheapest edge (under the constant elevation function), and cross using this function. Thus, the optimal path in this case can be restricted to use the grid edges, see Figure 4.4. The case that the path does the same thing in a row can be handled similarly.
@@ 4.5.4 Discussion of the two algorithms
Experiments show that the upper bound computed by the algorithm described above is a good approximation to the optimal solution (even with no refinement). This holds because some the VE graph edges (specifically, left-to-right, and bottom-to-top) realizes the minimum lines in the cell of the grid, thus at least one of the two integrals is minimized if the solution follow this segment, see Observation A.2. Thus, the paths suggested by the VE graph are already “cheap”, at least locally.
@@ 4.6.1 Fast extractor
Given a curve $\pi=p_{1}\ldots p_{n}$ with $n$ vertices, and consider any segment $\tau$. One can compute in linear time a monotone morphing $\mathcalb{m}$ between $\pi$ and $\tau$, such that $\mathsf{d}_{\mathcal{F}}\left(\pi,\tau\right)\leq D(\pi,\tau)\leq 3\mathsf{d}_{\mathcal{F}}\left(\pi,\tau\right)$, where $D(\pi,\tau)=\omega\left(\mathcalb{m}\right)$.
@@
Consider first the “silly” morphing $\mathcalb{m}^{\prime}$, that matches every vertex of $\pi$ to its nearest neighbor on $\tau$ (in between, it interpolates linearly). Clearly, this is a weak morphing between $\pi$ and $\tau$, but it is also clearly optimal. Consider the monotone version $\mathcalb{m}$ of $\mathcalb{m}^{\prime}$ that one gets by never moving back along $s$, and let $\ell$ be its width.
If $\mathcalb{m}^{\prime}$ is monotone, then $\mathcalb{m}=\mathcalb{m}^{\prime}$ and $\ell=\mathsf{d}_{\mathcal{F}}\left(\pi,\tau\right)$, and we are done. Otherwise, consider a point $q\in\tau$ such that the morphing goes over it several times, and let $\sigma=\pi\left\langle{q}\right\rangle$ be the portion of $\pi$ between the first time, and the last time, the morphing goes over $q$. Let $q_{s}$ be the leftmost point on $\tau$ matched by $\mathcalb{m}^{\prime}$ during the motion on $\pi\left\langle{q}\right\rangle$. Let $t=(q+q_{s})/2$. The optimal (monotone) morphing $\mathcalb{m}^{\star}$ matches $\sigma$ to segment $\nu\subseteq\tau$. Assume $q$ was chosen such that the length of $\nu$ is maximized among all such segments. If $t\in\nu$, then
@@
Indeed, the starting endpoint $\pi\left\langle{q}\right\rangle$ is matched to a point that appears in $\tau$ before $t$. Similar argument implies the same inequality if $t$ appears or after $\nu$ along $\tau$, as an easy case analysis applies. On the hand, since $\nu$ is of maximal length, we have that
$\omega\left(\mathcalb{m}\right)\leq\sqrt{\omega\left(\mathcalb{m}^{\prime}\right)^{2}+\left\|{\nu}\right\|^{2}}\leq\sqrt{2}\max(\omega\left(\mathcalb{m}^{\prime}\right),\left\|{\nu}\right\|)\leq 2\sqrt{2}\max(\omega\left(\mathcalb{m}^{\prime}\right),\left\|{\nu}\right\|/2)\leq 3\omega\left(\mathcalb{m}^{\star}\right).$
@@
The idea is to build a hierarchical representation of the curve. So let the input curve be $\pi=p_{1}p_{2}\ldots p_{n}$ – the output will be an array $A[1\ldots n]$ of real numbers. The roughest approximation for $\pi$ is the spine $p_{1}p_{n}$. If we want a finer approximation, the natural vertex to add is $p_{\alpha(1,n)}$, where $k=\alpha(i,j)=\left\lfloor{(i+j)/2}\right\rfloor$, which yields the curve $p_{1}p_{k}p_{n}$. The recursive algorithm ${{\color[rgb]{0.59,0,0.66}\definecolor[named]{pgfstrokecolor}{rgb}{0.59,0,0.66}\pgfsys@color@cmyk@stroke{0.07}{0.90}{0}{0.34}\pgfsys@color@cmyk@fill{0.07}{0.90}{0}{0.34}\texttt{\bf{comp{}Profile}}}}(1,n)$, would use the fast approximation algorithm of Lemma 4.24, the algorithms computes
$A[k]=\max(D(\pi[1,k]),D(\pi[k,n))].$
@@
Given a parameter $w$, the algorithm performs a recursive traversal of the curve, as described above. If the traversal algorithm arrives to an interval $\left\llbracket i:j\right\rrbracket$, with $k=\alpha(i,j)$, such that $A[k]>w$, then the algorithm adds $p_{k}$ to the simplified curve, after extracting the simplification recursively on the range $\left\llbracket i:k\right\rrbracket$, and before extracting the simplification $\left\llbracket k:j\right\rrbracket$. If $A[k]\leq w$, then the algorithm just adds $p_{k}$ to the simplification, without performing the recursive calls.
@@
Given a polygonal curve $\pi$ with $n$ vertices, the algorithm compProfile preprocess it, in $O(n\log n)$ time, such that given a parameter $w\geq 0$, the query algorithm extract computes a subcurve $\overline{\pi}$ (induced by a subset of the vertices of $\pi$), such that $\mathsf{d}_{\mathcal{F}}\left(\pi,\overline{\pi}\right)\leq w$. The extraction algorithm works in $O(\left|{V\left(\overline{\pi}\right)}\right|)$.
@@
A better simplification algorithm, that yields fewer vertices than the $\delta$-simplification used above, and the fast extractor used above, is to do a greedy approximation as suggested by Aronov et al. ahkww-fdcr-06. Let $\pi=p_{1}\ldots p_{n}$. The algorithm initially start at the vertex $p_{1}$. Assume that it is currently at the vertex $p_{j}$. The algorithm computes the first $k$, such that $\mathsf{d}_{\mathcal{F}}\left(\pi[p_{j},p_{k}],p_{j}p_{k},\leq\right)\delta$. One can use the algorithm Lemma 4.24 coupled with exponential and binary search to compute $k$ in time $O((k-j+1)\log(k-j+1))$ (of course, here the guarantee is somewhat weaker). This algorithm yields quite a good approximation in practice with fewer vertices than the $\delta$-simplification.

### v2

@@ 4.4.1 Morphing sensitive simplification
For any $\psi>1$, a Fréchet morphing $\mathcalb{m}$ of $\pi$ and $\sigma$ is a $\psi$-approximate morphing if $\mathsf{d}_{\mathcal{F}}\left(\pi,\sigma\right)\leq\omega\left(\mathcalb{m}\right)\leq\psi\mathsf{d}_{\mathcal{F}}\left(\pi,\sigma\right).$ Thus $\mathcalb{l}=\omega\left(\mathcalb{m}\right)/\psi$ is a lower bound on the Fréchet distance between $\pi$ and $\sigma$.
@@
For a morphing $\mathcalb{m}\in\mathcal{M}_{\pi,\sigma}$, a lower bound $\mathcalb{l}\leq\mathsf{d}_{\mathcal{F}}\left(\pi,\sigma\right)$, and a point $p=(x,y)\in\mathcalb{m}$, the slack at $p$ is $\Delta\left(p\right)=\max(\mathcalb{l}-\mathcalb{e}\left(p\right),0)$. A point $p\in\mathcalb{m}$ is tight if $\Delta\left(p\right)=0$.
Consider an optimal (say retractable) Fréchet morphing $\mathcalb{m}$ between $\pi$ and $\sigma$. If the lowerbound is the Fréchet distance, then we expect only the point in the morphing realizing the bottleneck to be tight. One can map the slack from the morphing back to the original curves. Specifically, for $v\in\pi\cup\sigma$, let $\max{m}(v)$ be the maximum length leash attached to $v$ during the morphing $\mathcalb{m}$ (this can correspond to the maximum elevation along a vertical or horizontal segment in $\mathcalb{m}$).
@@
For a vertex $v\in\pi\cup\sigma$, the slack of $v$ is $\Delta\left(v\right)=\max(\mathcalb{l}-\max\nolimits{m}(v),0).$
@@ 4.4.2 Computing a lower-bound on the Fréchet distance
$\leq\left\|{\overline{\pi}(\mathcalb{m}_{\pi}(x)\bigr{)}-\overline{\sigma}\bigl{(}\mathcalb{m}_{\sigma}(y)\bigr{)}}\right\|-\left\|{\pi(x)-\overline{\pi}(\mathcalb{m}_{\pi}(x)\bigr{)}}\right\|-\left\|{\sigma(y)-\overline{\sigma}(\mathcalb{m}_{\sigma}(y)\bigr{)}}\right\|\leq\mathcalb{e}\left(x,y\right).$
@@ 4.4.3 Computing the exact Fréchet distance
Let $\pi$ and $\sigma$ be two curves, and let $\mathcalb{m}$ be a $\psi$-approximate morphing between them, for some $\psi>1$. Let $\tau\geq 1$ be some constant, and consider $(\mathcalb{m},\tau)$-sensitive simplifications $\overline{\pi},\overline{\sigma}$, of $\pi,\sigma$, respectively. Let $\mathcalb{m}_{2}:\overline{\pi}\rightarrow\overline{\sigma}$ be the optimal Fréchet morphing between $\overline{\pi}$ and $\overline{\sigma}$, and let $\mathcalb{m}_{1}:\pi\rightarrow\overline{\pi}$, and $\mathcalb{m}_{3}:\overline{\sigma}\rightarrow\sigma$, be the natural morphings associated with the simplifications. This gives rise to a natural morphing $\mathcalb{m}^{\prime}=\mathcalb{m}_{3}\circ\mathcalb{m}_{2}\circ\mathcalb{m}_{1}$ from $\pi$ to $\sigma$. If $\omega\left(\mathcalb{m}^{\prime}\right)={\overline{\mathsf{d}_{\mathcal{F}}}\left(\overline{\pi},\overline{\sigma}\right)}$, then $\mathcalb{m}^{\prime}$ realizes the optimal Fréchet distance between $\pi$ and $\sigma$. where ${\overline{\mathsf{d}_{\mathcal{F}}}\left(\overline{\pi},\overline{\sigma}\right)}$ is the Fréchet distance between $\overline{\pi}$ and $\overline{\sigma}$ under the modified elevation function $\overline{\mathcalb{e}}$.
@@ 4.5.2 The sweep distance
$F(x)=\int f(x)\mathop{\mathrm{d}}\!x=\left(\frac{c}{2\sqrt{a}}-\frac{b^{2}}{8a^{3/2}}\right)\mathop{\mathrm{asinh}}\frac{2ax+b}{\sqrt{4ac-b^{2}}}+\left(\frac{x}{2}+\frac{b}{4a}\right)f(x).$
In particular, $\mathrm{price}\left(\tau\right)=F(\left\|{\tau}\right\|)-F(0)$ is the price of the CDTW charged to $\tau$. We repeat the same argument, but now the morphing is interpreted being for a point moving on $\tau^{\prime}$. This would yield a similar indefinite integral $G(\cdot)$, and the price of $\tau^{\prime}$ is $\mathrm{price}\left(\tau^{\prime}\right)=G(\left\|{\tau^{\prime}}\right\|)-G(0)$.
In the free space diagram, the morphing between $\tau$ and $\tau^{\prime}$ corresponds to a segment that lies in a cell, and $\mathrm{price}\left(\tau\right)+\mathrm{price}\left(\tau^{\prime}\right)$ is the price the sweep distance assigns this segment in the free space diagram.
@@
The above gives us a new pricing of the edges of the VE graph (see Definition 3.2), as every edge is a segment in the free space diagram. We can then use Dijkstra to compute the shortest path in the VE graph. Naturally, the resulting path is not monotone, but monotonicity can be easily achieved by introducing middle points, as described in Remark 4.6. We then recompute the shortest path getting a morphing, and repeat the refinement step if the morphing is not yet monotone. Usually, one round of refinement seems to suffice. We output the computed distance (for the associated monotone morphing). We refer to this quantity, as the sweep distance between the two original curves, denoted by $\mathsf{d}_{SW}\left(\pi,\sigma\right)$.
@@ 4.5.3 Computing a lower bound on the CDTW distance using the sweep distance
Consider computing the sweep distance for this “elevation” function. Since the sweep distance is decomposed along its $x$ and $y$ components, and the function is a constant inside a grid cell, one can safely assume the optimal morphing is axis aligned. The only remaining possibility is that the optimal morphing enters a column on a left edge, climbs in a stairway, and leaves through an edge on the right. Let $e_{1},e_{2},\ldots,e_{k}$ be the horizontal edges this path intersects, and observe that one can always modify this path, to move vertically to the cheapest edge (under the constant elevation function), and cross using this function. Thus, the optimal path in this case can be restricted to use the grid edges, see Figure 4.4. The case that the path does the same thing in a row can be handled similarly.
@@ 4.5.4 Discussion of the two algorithms
Experiments show that the upper bound computed by the algorithm described above is a good approximation to the optimal solution (even with no refinement). This holds because some of the VE graph edges (specifically, left-to-right, and bottom-to-top) realize the minimum lines in the cell of the grid, thus at least one of the two integrals is minimized if the solution follow this segment, see Observation A.2. Thus, the paths suggested by the VE graph are already “cheap”, at least locally.
@@ 4.6.1 Fast extractor
For a closed set $X\subseteq\mathbb{R}^{d}$ and a point $p\in\mathbb{R}^{d}$, let $\mathcalb{d}\left(p,X\right)=\min_{t\in X}\left\|{pt}\right\|$ be the distance of $p$ to $X$. Similarly, let $\Downarrow_{X}\!\!\left(p\right)=\arg\min_{t\in X}\left\|{pt}\right\|$ be the nearest point of $X$ to $p$. The projection of a set $Y$ to a closed set $X\subseteq\mathbb{R}^{d}$, is the set $\Downarrow_{X}\!\!\left(Y\right)=\left\{\Downarrow_{X}\!\!\left(p\right)\;\middle|\;p\in Y\right\}$.
@@
Given a (directed) curve $\pi=p_{1}\ldots p_{n}$ with $n$ vertices, consider any (directed) segment $\tau=zz^{\prime}$, such that $\Downarrow_{\tau}\!\!\left(p_{1}\right)=z$ and $\Downarrow_{\tau}\!\!\left(p_{n}\right)=z^{\prime}$. One can compute in linear time a monotone morphing $\mathcalb{m}$, and its width $\omega\left(\mathcalb{m}\right)$, between $\pi$ and $\tau$, such that $\mathsf{d}_{\mathcal{F}}\left(\pi,\tau\right)\leq\omega\left(\mathcalb{m}\right)\leq 3\mathsf{d}_{\mathcal{F}}\left(\pi,\tau\right)$.
@@
Consider first the morphing $\mathcalb{m}^{\prime}$, that maps each point of $\pi$ to its nearest point on $\tau$. Clearly, $\omega\left(\mathcalb{m}^{\prime}\right)\leq\omega\left(\mathcalb{m}^{\star}\right)$, where $\mathcalb{m}^{*}$ is the optimal (monotone) morphing between $\pi$ and $\tau$. Consider the monotone version $\mathcalb{m}$ of $\mathcalb{m}^{\prime}$ that one gets from $\mathcalb{m}^{\prime}$ by never moving back along $\tau$. If $\mathcalb{m}^{\prime}$ is monotone, then $\mathcalb{m}=\mathcalb{m}^{\prime}$ and $\omega\left(\mathcalb{m}\right)=\mathsf{d}_{\mathcal{F}}\left(\pi,\tau\right)$, and we are done.
Refer to caption Refer to caption
Figure 4.5: Illustration of the proof of Lemma 4.24.
The curve $\pi$ is uniformly parameterized on the interval $I=[0,\left\|{\pi}\right\|]$, see Definition 1.1. The morphing $\mathcalb{m}^{\prime}$ matches $\pi(t)$ with the point $\Downarrow_{\tau}\!\!\left(\pi(t)\right)$. Let $f(t)=\left\|{\Downarrow_{\tau}\!\!\left(\pi(t)\right)-z}\right\|$ be the normalized parameterization of this projection point on $\tau$. Thus, for all $t$, $f(t)\in[0,\left\|{\tau}\right\|]$, $f(0)=0$, and $f(\left\|{\pi}\right\|)=\left\|{\tau}\right\|$. The max version of $f$ is $F(t)=\max_{x:[0,t]}f(x)$, and the morphing $\mathcalb{m}$, for all $t$, maps $\pi(t)$ to $\tau(F(t))$. A maximal interval $J=[\alpha,\beta]\subseteq[0,\left\|{\pi}\right\|]$, with $\alpha<\beta$, such that $F(J)$ is a single value, is a plateau. For a plateau the morphing $\mathcalb{m}$ matches all the points of $\pi(J)$ to the point $q=\tau(F(J))$. A point $x\in[0,\left\|{\pi}\right\|]$ not contained in any plateau, is matched by $\mathcalb{m}$, to the same point it was matched by $\mathcalb{m}^{\prime}$, that is $\tau(f(x))=\Downarrow_{\tau}\!\!\left(\pi(x)\right)$.
The only places where the leash used by $\mathcalb{m}$ might be larger than $\mathcalb{m}^{\star}$ is in the plateaus. So consider such a plateau $J=[\alpha,\beta]$, with $\alpha<\beta$. Let $\nu=\left\{\tau(f(t))\;\middle|\;t\in J\right\}$ be the portion of $\tau$ that is being used by $\mathcalb{m}^{\prime}$ for points of $\sigma=\pi(J)$. The start of this plateau in $\pi$ is $s=\pi(\alpha)$, and the end is $t=\pi(\beta)$. The left endpoint of $\nu\subseteq\tau$ is the projection of some point $u=\pi(\gamma)$, with $\gamma\in(\alpha,\beta)$, such that $f(J)=[f(\gamma),f(\beta)]$. Let $q_{L}=\tau(f(\gamma))$ and $q=\tau(f(\beta))$ be the right and left endpoints of $\nu$, respectively, see Figure 4.5. Finally, let $r$ be the middle point of $\nu$.
If the optimal (monotone) morphing $\mathcalb{m}^{\star}$ matches $s$ to a point $s^{\prime}\in\tau$ that appears before $r$ on $\tau$, then $\omega\left(\mathcalb{m}^{\star}\right)\geq\left\|{ss^{\prime}}\right\|\geq\left\|{\nu}\right\|/2$, see Figure 4.5 (left). Otherwise, $\mathcalb{m}^{\star}$ maps $u$ to a point $u^{\prime}\in\tau$ that appears after $r$ on $\tau$. But then, $\omega\left(\mathcalb{m}^{\star}\right)\geq\left\|{uu^{\prime}}\right\|\geq\left\|{\nu}\right\|/2$, see Figure 4.5 (right). We conclude that
@@
As a reminder, $\mathcalb{m}$ matches all the points of $\sigma=\pi(J)$ to $q$. Let $\ell=\max_{p\in\sigma}\mathcalb{d}\left(p,\nu\right)$ be the maximum leash length deployed by $\mathcalb{m}^{\prime}$ for any point of $\sigma$. We have that the width of $\mathcalb{m}$ restricted to $\sigma$ is
$\ell_{J}$ $=\max_{p\in\sigma}\left\|{pq}\right\|\leq\sqrt{\ell^{2}+\left\|{\nu}\right\|^{2}}\leq\sqrt{\omega\left(\mathcalb{m}^{\prime}\right)^{2}+\left\|{\nu}\right\|^{2}}\leq\sqrt{2}\max(\omega\left(\mathcalb{m}^{\prime}\right),\left\|{\nu}\right\|)$
$\leq 2\sqrt{2}\max(\omega\left(\mathcalb{m}^{\prime}\right),\left\|{\nu}\right\|/2)\leq 3\omega\left(\mathcalb{m}^{\star}\right).$
Now, $\omega\left(\mathcalb{m}\right)$ is the maximum of $\omega\left(\mathcalb{m}^{\prime}\right)$ and the maximum of $\ell_{J}$, over all plateaus $J$, which implies the claim.
@@
The idea is to build a hierarchical representation of the curve. So let the input curve be $\pi=p_{1}p_{2}\ldots p_{n}$ – the output will be an array $A[1\ldots n]$ of real numbers. The roughest approximation for $\pi$ is the spine $p_{1}p_{n}$. If we want a finer approximation, the natural vertex to add is $p_{\alpha(1,n)}$, where $k=\alpha(i,j)=\left\lfloor{(i+j)/2}\right\rfloor$, which yields the curve $p_{1}p_{k}p_{n}$. The recursive algorithm ${{\color[rgb]{0.59,0,0.66}\definecolor[named]{pgfstrokecolor}{rgb}{0.59,0,0.66}\pgfsys@color@cmyk@stroke{0.07}{0.90}{0}{0.34}\pgfsys@color@cmyk@fill{0.07}{0.90}{0}{0.34}\texttt{\bf{comp{}Profile}}}}(1,n)$ uses the fast approximation algorithm of Lemma 4.24 to compute
$A[k]=\max(D(\pi[1,k]),D(\pi[k,n])].$
@@
Given a parameter $w$, the algorithm performs a recursive traversal of the curve, as described above. If the traversal algorithm arrives to an interval $\left\llbracket i:j\right\rrbracket$, with $k=\alpha(i,j)$, such that $A[k]>w$, then the algorithm adds $p_{k}$ to the simplified curve, after extracting the simplification recursively on the range $\left\llbracket i:k\right\rrbracket$, and before extracting the simplification recursively on $\left\llbracket k:j\right\rrbracket$. If $A[k]\leq w$, then the algorithm just adds $p_{k}$ to the simplification, without performing the recursive calls.
@@
Given a polygonal curve $\pi$ with $n$ vertices, the algorithm compProfile preprocess it, in $O(n\log n)$ time, such that given a parameter $w\geq 0$, the query algorithm extract computes a subcurve $\overline{\pi}$ (induced by a subset of the vertices of $\pi$), such that $\mathsf{d}_{\mathcal{F}}\left(\pi,\overline{\pi}\right)\leq w$. The extraction algorithm works in $O(\left|{V\left(\overline{\pi}\right)}\right|)$ time.
@@
A better simplification algorithm, that yields fewer vertices than the $\delta$-simplification used above, and the fast extractor used above, is to do a greedy approximation as suggested by Aronov et al. ahkww-fdcr-06. Let $\pi=p_{1}\ldots p_{n}$. The algorithm initially starts at the vertex $p_{1}$. Assume that it is currently at the vertex $p_{j}$. The algorithm then sets the next vertex in the simplification to be $p_{k}$, where $k$ is the smallest index greater than $j$ such that $\mathsf{d}_{\mathcal{F}}\left(\pi[p_{j},p_{k}],p_{j}p_{k}\right)\leq\delta$ and $\mathsf{d}_{\mathcal{F}}\left(\pi[p_{j},p_{k+1}],p_{j}p_{k+1}\right)>\delta$ (or $k=n$ if the end of the curve is reached). One can use the algorithm Lemma 4.24 coupled with exponential and binary search to compute $k$ in time $O((k-j+1)\log(k-j+1))$ (of course, here the guarantee is somewhat weaker). This algorithm yields quite a good approximation in practice with fewer vertices than the $\delta$-simplification.